# TransDiffuser: End-to-end Trajectory Generation with Decorrelated Multi-modal Representation for Autonomous Driving

### v1

@@ 5 $\mathcal{L}_{diff}\leftarrow\text{MSE}(pred,noise)$ ;
During the inference process, the model leverages this learned denoising process to generate a pool of diverse candidate trajectories starting from pure noise. In our implementation, we generate 30 candidates considering efficiency. Subsequently, a rejection sampling [9] strategy is employed to filter out dynamically infeasible or unsuitable trajectories. This contrasts with methods like GoalFlow [10] generating 128/256 candidates or TrajHF [9] using 100 candidates before selection or filtering. This reflects the inference efficiency of our method.
3.4 Multi-modal Representation Decorrelation
The final denoised action is decoded by the action decoder as illustrated in Figure 2. We believe that the final performance is largely determined by the quality of the learned multi-modal fused representations. To tackle this information bottleneck [42], we apply multi-modal representation optimization objective on the fused multi-modal representations, which is illustrated in Figure 2. In detail, the input multi-modal representation matrix $\mathbf{M}$ (shown in Figure 1) with shape $\mathbf{B}\times\mathbf{D}$, where $\mathbf{B}$ and $\mathbf{D}$ denote the batchsize and the reshaped multi-modal representation dimension of each sample. Detailedly, in our implementation, $\mathbf{M}$ is reshaped from $B\times 8\times 256$ to $B\times 2048$. We first normalize multi-modal representation matrix $\mathbf{M}$ and then compute its correlation matrix corr. This penalty $\mathcal{L}_{reg}$ is designed to decrease non-diagonal entries of the multi-modal correlation matrix corr.
Figure 2: Illustrations on the multi-modal representation decorrelation process. (d) shows top 100 largest singular values of the correlation matrices. We observe the proposed decorrelation mechanism helps to better utilize the representation space before action decoding (shown in Figure 1).
Intuitively, $\mathcal{L}_{reg}$ aims to regularize the multi-modal correlation matrix (i.e. Figure 2) to be similar to the diagonal form. This can be achieved by eliminating the unnecessary interactions of different dimensions of representation to increase the similarity (illustrated as $sim(\cdot)$) to identity matrix or diagonal matrix (i.e. Figure 2(c)). As the figure shows, the derived $\mathcal{L}_{rep}$ is a batch-level regularization objective which optimizes the multi-modal representations in each training batch. Since the singular values of the covariance matrix corr provide a comprehensive characterization [31] of the distribution of multi-modal representations M, we visualize the largest singular values shown in Figure 2(d). We also provide the batch-level sensitivity study in Section 4.3.
Input: Reshaped Multi-modal representation matrix $\mathbf{M}$
@@ 3.4 Multi-modal Representation Decorrelation
2 $\mathbf{M}\leftarrow\frac{\mathbf{M}}{\sqrt{\epsilon+\sigma_{\mathbf{M}}}}$ ; //Normalization. $\epsilon=1e^{-8}$ is used to avoid zero denominator.
3 $\mathbf{corr}\leftarrow\mathbf{M}^{T}\cdot\mathbf{M}$ ; //Compute the correlation matrix.
4 $\hat{corr}\leftarrow\text{remove\_diagonal\_elements}(\mathbf{corr})$ ; //Extract the non-diagonal elements.
5 $\mathcal{L}_{rep}\leftarrow(\hat{corr}^{2}.mean())/\mathbf{B}$ ;
@@
During the training process, we combine the supervised $\mathcal{L}_{diff}$ and the multi-modal representation decorrelation loss with the trade-off coefficient $\beta$ to form the overall optimizing objective:
@@
4 Experiments
4.1 Experimental Setup
@@
Dataset. We utilize the well-established planning-oriented NAVSIM dataset [43] using non-reactive simulation and closed-loop evaluations. It builds on the existing OpenScene [44] dataset, a compact version of nuPlan [45] sampled at 2 Hz. Each sample contains camera images from 8 perspectives, fused LiDAR data from 5 sensors, ego status, and annotations for the map and objects. It has two parts: Nav-train and Nav-test, which respectively contain 1192 and 136 scenarios for training/validation and testing, and we only use the training split for training and validation set to guide model selection.
@@
Baseline. In this work, we compare the proposed framework against counterparts from three groups discussed in Section 2. Following  [43, 10], we additionally refer to Constant Velocity and Ego Status MLP as the lower bound for comparison. Constant Velocity model assumes constant speed from the current timestamp for forward movement. Ego Status MLP serves as a bind driving agent, which leverages a MLP for trajectory prediction given only the ego vehicle status.
@@
Implementation Details. Our code is built on the PyTorch lightening framework [46] and official NAVSIM toolkit [43]. We take 10 timesteps during denoising process for both training and inference. The trade-off coefficient $\beta$ in Eq. 4 is fixed to $0.02$. The learning rate is set to $41e-$ 4, and global batchsize is $256$, distributed across $4$ NVIDIA® H20 GPUs. Adam optimizer [47] is adopted with OneCycle scheduler. As the task requires, the model outputs $8$ waypoints spanning $4$ seconds. Total training lasts for $120$ epochs.
@@
$PDM_{score}=NC\times DAC\times TTC\times\frac{(5\times DDC+2\times C+5\times EP)}{12},$ (5)
@@
where $NC$, $DAC$, $TTC$, $C$, $EP$, $DDC$ correspond to the No at-fault Collisions, Drivable Area Compliance, Time to Collision, Comfort, Ego Progress, and Driving Direction Compliance. Note that DDC is exempted from calculation due to the practical constraints of the NAVSIM toolkit [^2] https://github.com/autonomousvision/navsim/issues/14. Following [16, 9], top-1 scoring predicted trajectory for each sample is used for PMDS evaluation. To measure the diversity of all generated candidate trajectories for each sample, we refer to [16] for the $\mathcal{D}$iversity metric, and further discuss in Section 4.2.
Table 2: Performance on Navtest with closed-loop metrics. L , V and V* denote the LiDAR input, vision input, and the video or historical image frame input, respectively.
Method Modality Metrics
$NC\uparrow$ $DAC\uparrow$ $EP\uparrow$ $TTC\uparrow$ $C\uparrow$ $PDMS\uparrow$
Constant Velocity [43] - 68.0 57.8 19.4 50.0 100.0 20.6
Ego Status MLP [43] - 93.0 77.3 62.8 83.6 100.0 65.6
Transfuser  [8] V+L 97.8 92.6 78.9 92.9 100.0 83.9
LTF [8] V 97.4 92.8 79.0 92.4 100.0 83.8
UniAD [1] V+L 97.8 91.9 78.8 92.9 100.0 83.4
PARA-Drive [19] V*+L 97.9 92.4 79.3 93.0 99.8 84.0
DiffusionDrive  [16] V+L 98.2 96.2 82.2 94.7 100.0 88.1
GoalFlow  [10] V+L 98.4 98.3 85.0 94.6 100.0 90.3
VADv2 [12] V+L 97.2 89.1 76.0 91.6 100.0 80.9
Hydra-MDP  [13] V+L 99.1 98.3 85.2 96.6 100.0 91.3
Hydra-NeXt  [15] V 98.1 97.7 81.8 94.6 100.0 88.6
Hydra-MDP++ (ResNet34)  [14] V* 97.6 96.0 80.4 93.1 100.0 86.6
WoTE [21] V+L 98.5 95.8 80.9 94.4 99.9 87.1
Hydra-MDP++ (V2-99)  [14] V* 98.6 98.6 85.7 95.1 100.0 91.0
Centaur [20] V+L 99.5 98.9 85.9 98.0 100.0 92.6
TrajHF [9] V+L 99.3 97.5 90.4 98.0 99.8 94.0
TransDiffuser (Ours) V+L 99.4 96.5 94.1 97.8 99.4 94.9
4.2 Main Experiments
@@
Benchmark Performance. As Table 2 indicates, TransDiffuser achieves the best performance on Nav-test. Performance of other models are cited from the original papers. Hydra-MDP++ provides a base version and a large version by scaling up the image encoder from ResNet-34 [49] to V2-99 [50]. The most obvious improvement over previous works is the Ego Progress metric, which can also be intuitively observed in Figure 3. We can also witness the potential of Diffusion based models. Note that we do not require any anchor distributions or anchor trajectories as prior, discussed in Section 2.
@@
Training Efficiency. Our training performs 120 epochs across four GPUs and consumes within 2 wall-clock hours. TransDiffuser model is about 251M parameters while the parameters of the perception encoder are frozen, thus only 62.8M parameters are trainable. Regarding inference-time efficiency, we discussed in Section 3.3.
Diversity Improvement. Following the diversity metric proposed and implemented in DiffusionDrive [16], as shown in Table 3, we additionally measure the averaged $\mathcal{D}iversity$ metric of the generated trajectories of each sample over the test dataset. This measures the mean Intersection over Union (mIoU) between each denoised generated trajectory and the union of all denoised candidate trajectories for each sample. From the experimental results, we observe the existence of Multi-modal representation decorrelation regularization correspondingly improves the diversity of generated candidate trajectories (from $65.90$ to $70.02$).
4.3 Ablation & Sensitivity Study
@@
Herein we study three key hyper-parameters: Diffusion timesteps $\mathbf{T}$ in Section 3.3, Batchsize $\mathbf{B}$ in Section 3.4 and $\beta$ in Eq.3. As Table 3 indicates, our method shows robustness over the selection of coefficient $\beta$, diffusion timesteps $\mathbf{T}$ and batchsize $\mathbf{B}$. We obverse when we scale up the diffusion timesteps, the $\mathcal{D}iversity$ metric generally increases at the cost of more computation and latency, while small diffusion timesteps can also achieve satisfactory performance with higher efficiency.
Table 3: Exprimental results for ablation and sensitivity study.
@@
$NC\uparrow$ $DAC\uparrow$ $EP\uparrow$ $TTC\uparrow$ $C\uparrow$ $PDMS\uparrow$ $\mathcal{D}iversity\uparrow$
Timestep $\mathbf{T}$ 5 98.5 94.4 92.7 95.8 99.8 92.4 65.35
10 99.4 96.5 94.1 97.8 99.4 94.9 70.02
20 99.0 96.0 94.6 96.7 99.3 94.3 87.99
Batchsize $\mathbf{B}$ 32 98.8 94.6 92.9 96.1 99.9 92.7 65.60
64 99.4 96.5 94.1 97.8 99.4 94.9 70.02
128 98.9 95.4 91.7 97.0 99.0 92.9 69.08
Coefficient $\beta$ 0 99.0 95.8 94.5 96.8 99.9 94.3 65.90
0.02 99.4 96.5 94.1 97.8 99.4 94.9 70.02
0.05 99.0 95.9 93.8 97.2 99.3 94.1 69.26
0.1 99.0 95.9 93.3 97.3 99.1 94.0 69.45
4.4 Qualitative Analysis

### v2

@@ 9 $\mathcal{L}_{diff}\leftarrow\text{MSE}(pred,noise)$ ;
During the inference process, the model leverages this learned denoising process to generate a pool of diverse candidate trajectories starting from pure Gaussian noise. In our implementation, we generate $N=30$ candidate trajectories considering the efficiency. Subsequently, rejection sampling strategy is employed to filter out dynamically infeasible or unsuitable trajectories following TrajHF [8]. sensitivity analysis on the candidate number is provided in Section IV-C.
Discussion: GoalFlow [9] typically generates 128 or 256 candidate trajectories, while TrajHF [8] uses 100 candidates prior to their respective selection or filtering steps. As analyzed in Section IV-D, we generate fewer candidates but still maintain effective trajectory quality, our method naturally embodies a degree of inference efficiency.
III-D Multi-modal Representation Decorrelation
The final denoised action is decoded by the action decoder as illustrated in Figure 3. We argue planning performance is determined by the quality of the learned multi-modal representations to certain degree. To tackle this information bottleneck [43], we apply the multi-modal representation optimization objective on the fused multi-modal representations, which is illustrated in Figure 3. In each training batch $B$, we first normalize multi-modal representation matrix $\mathbf{M}$ and then compute its correlation matrix corr. This penalty $\mathcal{L}_{reg}$ is designed to decrease non-diagonal entries of the multi-modal correlation matrix corr.
Figure 3: Illustrations on the multi-modal representation decorrelation process. (d) shows top 100 largest singular values of the correlation matrix. We observe the decorrelation mechanism helps to better utilize the representation space before action decoding (shown in Figure 2).
Intuitively, $\mathcal{L}_{reg}$ aims to regularize the multi-modal correlation matrix (i.e. Figure 3) to be similar to the diagonal form. This can be achieved by eliminating the unnecessary interactions of different dimensions of representation to increase the similarity (illustrated as $sim(\cdot)$) to identity matrix or diagonal matrix (i.e. Figure 3(c)). As the figure shows, the derived $\mathcal{L}_{rep}$ is a batch-level regularization objective which optimizes the multi-modal representations in each training batch. Since the singular values of the covariance matrix corr provide a comprehensive characterization [33] of the distribution of multi-modal representations M, we visualize the largest singular values shown in Figure 3(d). We also provide the batch-level sensitivity study in Section IV-C.
Input: Reshaped multi-modal representation $\mathbf{M}$
@@ III-D Multi-modal Representation Decorrelation
2 //Normalization. $\epsilon=1e^{-8}$ is used to avoid zero denominator.
3 $\mathbf{M}\leftarrow\frac{\mathbf{M}}{\sqrt{\epsilon+\sigma_{\mathbf{M}}}}$ ;
4 //Compute the correlation matrix.
5 $\mathbf{corr}\leftarrow\mathbf{M}^{T}\cdot\mathbf{M}$ ;
6 //Extract the non-diagonal elements.
7 $\hat{corr}\leftarrow\text{remove\_diagonal\_elements}(\mathbf{corr})$ ;
8 $\mathcal{L}_{rep}\leftarrow(\hat{corr}^{2}.mean())/|\mathbf{B}|$ ;
@@ 7 $\hat{corr}\leftarrow\text{remove\_diagonal\_elements}(\mathbf{corr})$ ;
We combine the supervised $\mathcal{L}_{diff}$ and the multi-modal representation decorrelation loss with the trade-off coefficient $\beta$ during training:
@@
Discussion: Herein we simply clarify this novelty. Though representation optimization is relatively widely studied in traditional machine learning tasks [44, 36, 35], it is rarely considered in generative models [45], especially for trajectory generation oriented at end-to-end autonomous driving. In the meantime, decorrelation related computation overhead (FLOPs) is less than 1‰ of the forward process and it benefits the generation diversity as shown in Figure 5.
III-E Quantification of Mode Collapse
To evaluate the multi-mode property of the generative trajectory models, we follow DiffusionDrive [14] to develop the quantification analysis. As analyzed in previous diffusion based studies [14, 9], different initialized random noises often lead to converged similar trajectories after the denoising process. To quantitatively analyze the phenomenon of mode collapse, we formulate the mode diversity score $\mathcal{D}$ based on the mean Intersection over Union (mIoU) between each denoised trajectory and the union of all denoised trajectories:
$\mathcal{D}=1-\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Area}\left(\tau_{i}\cap\bigcup_{j=1}^{N}\tau_{j}\right)}{\text{Area}\left(\tau_{i}\cup\bigcup_{j=1}^{N}\tau_{j}\right)}$ (5)
where $\tau_{i}$ represents the ithi-th denoised trajectory, $N$ is the total number of sampled trajectories and $\tau_{j}$ is the union of all denoised trajectories, as introduced in Section III-C.
IV Experiments
IV-A Experimental Setup
@@
Dataset. Following previous state-of-the-art works [14, 9, 8], We utilize the well-established planning-oriented NAVSIM dataset [46] using non-reactive simulation and closed-loop evaluations. It builds on the existing OpenScene [47] dataset, a compact version of nuPlan [48] sampled at 2 Hz. Each sample contains camera images from 8 perspectives, fused LiDAR data from 5 sensors, ego status, and annotations for the map and objects. It has two parts: Nav-train and Nav-test, which respectively contain 1192 and 136 scenarios for training/validation and testing, and we only use the training split for training and validation set to guide model selection.
@@
Baseline. We compare the proposed framework against counterparts from three groups discussed in Section II. Following  [46, 9], we additionally refer to Constant Velocity and Ego Status MLP as the lower bound for comparison. Constant Velocity model assumes constant speed from the current timestamp for forward movement. Ego Status MLP serves as a bind driving agent, which leverages a MLP for trajectory prediction given only the ego vehicle status.
TABLE II: Performance on NAVSIM. L, V and V* denote LiDAR input, vision input, and video or historical image input.
Method Modality Image Encoder Metrics
$NC\uparrow$ $DAC\uparrow$ $EP\uparrow$ $TTC\uparrow$ $Comfort\uparrow$ $\textbf{PDMS}\uparrow$
Constant Velocity [46] - - 68.0 57.8 19.4 50.0 100 20.6
Ego Status MLP [46] - - 93.0 77.3 62.8 83.6 100 65.6
Transfuser  [7] V+L ResNet-34 97.8 92.6 78.9 92.9 100 83.9
LTF [7] V ResNet-34 97.4 92.8 79.0 92.4 100 83.8
UniAD [5] V+L ResNet-34 97.8 91.9 78.8 92.9 100 83.4
PARA-Drive [17] V*+L ResNet-34 97.9 92.4 79.3 93.0 99.8 84.0
DiffusionDrive  [14] V+L ResNet-34 98.2 96.2 82.2 94.7 100 88.1
GoalFlow  [9] V+L ResNet-34 98.4 98.3 85.0 94.6 100 90.3
VADv2 [18] V+L ResNet-34 97.2 89.1 76.0 91.6 100 80.9
Hydra-MDP  [11] V+L ResNet-34 99.1 98.3 85.2 96.6 100 91.3
Hydra-NeXt  [13] V ResNet-34 98.1 97.7 81.8 94.6 100 88.6
Hydra-MDP++ (Base)  [12] V* ResNet-34 97.6 96.0 80.4 93.1 100 86.6
WoTE [20] V+L ResNet-34 98.5 95.8 80.9 94.4 99.9 87.1
Hydra-MDP++ (Large)  [12] V* V2-99 98.6 98.6 85.7 95.1 100 91.0
R2SE [21] V BEVFormer 99.0 97.9 86.8 96.4 100 91.6
Centaur [19] V+L ResNet-34 99.5 98.9 85.9 98.0 100 92.6
TrajHF [8] V+L ViT 99.3 97.5 90.4 98.0 99.8 94.0
DIVER [23] V+L ResNet-34 98.5 96.5 82.6 94.9 100 88.3
TransDiffuser (Ours) V+L ResNet-34 99.4 96.5 94.1 97.8 99.4 94.9
@@
Implementation Details. Our code is built on the PyTorch lightening framework [49] and official NAVSIM toolkit [46]. We take 10 timesteps during denoising process for both training and inference. The trade-off coefficient $\beta$ in Eq. 4 is fixed to $0.02$. The learning rate is set to 1e41e-4, and global batchsize is $256$, distributed across $4$ NVIDIA® H20 GPUs. Adam optimizer [50] is adopted with OneCycle scheduler. As the task requires, the model outputs $8$ waypoints spanning $4$ seconds. Total training lasts for $120$ epochs.
@@
$PDMS$ $=NC\times DAC\times TTC\times$
$\quad\frac{(5\times DDC+2\times C+5\times EP)}{12},$ (6)
@@
where the sub-metrics $NC$ (No at-fault Collision), $DAC$ (Drivable Area Compliance), $EP$ (Ego Progress), $Comfort$, DDCDDC (Driving Direction Compliance), and $TTC$ (Time-to-Collision), each represented as a percentage, are composed into a single score [46]. $DDC$ is exempted from calculation due to the practical constraints of the NAVSIM toolkit [^2] https://github.com/autonomousvision/navsim/issues/14. Following [14, 8], the top-1 scoring predicted trajectory for each sample is used for evaluation. To measure the diversity of all generated candidate trajectories for each sample, we refer to [14] for Diversity metric $\mathcal{D}$.
IV-B Main Experiments
@@
Benchmark Performance. As Table II indicates, TransDiffuser achieves the best performance on Nav-test against baseline methods like DiffusionDrive and GoalFlow. Performance of other models are cited from the original papers. Hydra-MDP++ provides a base version and a large version by scaling up the image encoder from ResNet-34 [52] to V2-99 [53]. The most obvious improvement of TransDiffuser lies in Ego Progress metric, which can also be intuitively observed in Figure 4 and 5. We also witness the potential of Diffusion based models. Note that different from previous state-of-the-art methods like GoalFlow and DiffusionDrive, we do not rely on any priors as discussed in Section II.
Diversity Improvement. Following the diversity metric in Section III-E, as shown in Table III and IV, we measure the $\mathcal{D}iversity$ metric over the test dataset. We observe the existence of multi-modal representation decorrelation regularization correspondingly improves the diversity of generated candidate trajectories (from $66$ to $70$) while there still exists a small gap to DiffusionDrive ($74$) which benefits from additional trajectory vocabulary to initialize the anchored distribution. In general, TransDiffuser can yield relatively diverse trajectories without any additional priors like anchored trajectories or scene priors. We conduct further analysis on how denoising steps and candidate number affect the diversity in Section IV-C and IV-D, which shows that TransDiffuser can achieves better diversity with more denoising timesteps, and it achieves better performance with 10/15 candidates than 20 candidates of DiffusionDrive.
@@
Training Efficiency. Our training performs 120 epochs across four GPUs and consumes within 2 wall-clock hours. TransDiffuser model is about 251M parameters while the parameters of the perception encoder are frozen, thus only 62.8M parameters are trainable. Regarding inference-time efficiency, we discussed in Section III-C.
IV-C Sensitivity Study
@@
Herein we study three key hyper-parameters: Diffusion timesteps $\mathbf{T}$ in Section III-C, Batchsize $\mathbf{B}$ in Section III-D and $\beta$ in Eq.3. As Table III indicates, our method shows robustness over the selection of coefficient $\beta$, diffusion timesteps $\mathbf{T}$ and batchsize $\mathbf{B}$. We obverse when we scale up the diffusion timesteps, the diversity metric $\mathcal{D}$ generally increases at the cost of more computation and latency, while small diffusion timesteps can also achieve satisfactory performance with higher efficiency and relatively lower diversity. Future works can consider more advanced policies like flow matching.
TABLE III: Experimental results for the sensitivity study.
@@
NCNC DACDAC EPEP TTCTTC ComfortComfort PDMS $\mathcal{D}$
Timestep $\mathbf{T}$ 5 98.5 94.4 92.7 95.8 99.8 92.4 65
10 99.4 96.5 94.1 97.8 99.4 94.9 70
20 99.0 96.0 94.6 96.7 99.3 94.3 88
Batchsize $\mathbf{B}$ 32 98.8 94.6 92.9 96.1 99.9 92.7 66
64 99.4 96.5 94.1 97.8 99.4 94.9 70
128 98.9 95.4 91.7 97.0 99.0 92.9 69
Coefficient $\beta$ 0 99.0 95.8 94.5 96.8 99.9 94.3 66
0.02 99.4 96.5 94.1 97.8 99.4 94.9 70
0.05 99.0 95.9 93.8 97.2 99.3 94.1 69
0.1 99.0 95.9 93.3 97.3 99.1 94.0 69
IV-D Exploration Experiments
Herein we discuss some key exploration regarding our TransDiffuser and multi-modal decorrelation optimization. Experimental results are shown in Table IV.
Candidiate number. The number of candidate trajectories is a key factor to performance. Previous methods like GoalFlow and TrajHF require no less than 100 candidate trajectories. With the assistance of anchored trajectory distribution and optimized generation policy, DiffusionDrive [14] exhibits satisfactory performance by generating 20 candidates. We decrease the candidate number to 10 and 15, our proposed TransDiffuser still shows its robustness on planning while fewer candidates yield the inferior diversity.
Inner or outer decorrelation. In our original design in Figure 2, the decorrelation is adopted on the outside of denoising decoder and before the action decoder. We also explore when we apply this decorrelation mechanism inside the denoising loop within the denoising decoder, and the results indicate this inner decorrelation shows slight performance and diversity decrease.
Holistic training. Scene encoder is frozen in TransDiffuser since we observe the slight performance decrease with the consistent diversity when we fully train the holistic model.
Generalization potential. We apply this decorrelation optimization on the auto-regressive based Transfuser, we find the planning performance gets improved.
TABLE IV: Performance on the exploration experiments.
Method Img. Encoder Anchor Candidate(s) PDMS $\mathcal{D}$
Transfuser ($\dagger$) ResNet-34 $\times$ 1 78.0 -
Transfuser ($\dagger$) + Decorr. ResNet-34 $\times$ 1 78.8 -
GoalFlow ResNet-34 $\checkmark$ 128/256 90.3 N/A
DiffusionDrive ResNet-34 $\checkmark$ 20 88.1 74
TrajHF ViT $\times$ 100 94.0 N/A
Ours ResNet-34 $\times$ 10 89.6 56
Ours ResNet-34 $\times$ 15 91.7 63
Ours ResNet-34 $\times$ 30 94.9 70
Ours (Fully-trained) ResNet-34 $\times$ 10 88.7 55
Ours (Fully-trained) ResNet-34 $\times$ 15 90.8 60
Ours (Fully-trained) ResNet-34 $\times$ 30 93.5 68
Ours (Inner.) ResNet-34 $\times$ 10 86.8 56
Ours (Inner.) ResNet-34 $\times$ 15 89.8 63
Ours (Inner.) ResNet-34 $\times$ 30 93.5 69
IV-E Qualitative Analysis